OpenStudy (anonymous):

Find the lesser of two consecutive integers with a sum greater than 16.

OpenStudy (jim766):

ok we have 2 numbers that we do not know...but we do know they are consecutive integers.... first think about consecutive integers, how much between each one?

OpenStudy (jim766):

1,2,3,4,5,6,7...

OpenStudy (anonymous):

1? @Jim766

OpenStudy (jim766):

right...right... so if the first number is going to be called x could we call the 2nd x + 1 ?

OpenStudy (anonymous):

is there more to this problem? because there are infinitely many answers...

OpenStudy (anonymous):

@dpaInc Yeah this is the entire problem, we're doing inequalities in Algebra II

OpenStudy (anonymous):

why an \(x\)? pick a number i pick 5 next number is 6 5+6=11 which in not greater than 16, so i try again try 7

OpenStudy (anonymous):

hmm... because this works 1000 and 1001.... consecutive integers, sum is greater than 16 and the smaller one is 1000

OpenStudy (anonymous):

\(7+8=15\) nope ok lets try 8 \(8+9=17\) hmmm

OpenStudy (anonymous):

@satellite73 ohh :)

OpenStudy (anonymous):

question asks for an inequality, so yes, there are an infinite number of answers to the equality

OpenStudy (anonymous):

i guess the number has to be greater than or equal to 8

OpenStudy (anonymous):

@satellite73 oh ok :) thank you